OpenStudy (anonymous):

Simplify the expression please! x sqrt8x ^3 + xsqrt2+4sqrt2x If the simplified expression is written in standard form, what is the LEADING COEFFCIENT? Please answer with a numerical answer.

OpenStudy (anonymous):

is this like this ? x sqrt(8x ^3) + xsqrt(2)+4sqrt(2x) ??

OpenStudy (anonymous):

Yes, sir.

OpenStudy (anonymous):

alright so lets use the property that sqrt(a*b) = sqrt(a)*sqrt(b) and lets break our formula up to get x*sqrt(8)*sqrt(x^3) +xsqrt(2) +4sqrt(2)*sqrt(x) and notice in the first term we got sqrt(8) we can also break that down to sqrt(4)*sqrt(2) = 2sqrt(2) and also tgat sqrt(x^3) = xsqrt(x) now lets plug everything back in to x*x*sqrt(x)*2sqrt(2) +x*sqrt(2) +4sqrt(2)*sqrt(x) which simplifies to 2(x^2)*sqrt(2)*sqrt(x) +x*sqrt(2) + 4sqrt(2)*sqrt(x) and the leading coefficient is 2^(3/2)

OpenStudy (anonymous):

Thank you so much!